# Early-Cycle Internal Impedance Enables ML-Based Battery Cycle Life Predictions Across Manufacturers

### v1

@@ 3 Machine Learning Approach
Nevertheless, to evaluate the generalization performance of the trained models, two distinct cross-validation strategies were employed, each focusing on different aspects of the dataset variability. A leave one triplicate out cross-validation strategy was adopted to assess the model’s ability to generalize across different operating conditions. The dataset is composed of cells grouped into triplicates, where each triplicate is a set of three cells cycled under identical conditions (i.e., a specific combination of manufacturer, temperature, and voltage range). In this approach, the model was trained on all but one triplicate, and the remaining triplicate was used for testing. This process was repeated until all triplicates were used as test sets, thus evaluating the model across a diverse range of operating conditions. By excluding entire triplicates during each fold, this method tests the model’s ability to predict cell behavior under previously unseen conditions, effectively simulating the real-world scenario where a battery management system must generalize to new operational environments. This strategy maximizes the utility of the dataset and provides a robust assessment of the model’s generalization capabilities across temperature, voltage, and manufacturer combinations.
In addition to operating conditions, cross-manufacturer validation was performed to assess the model’s ability to generalize across different cell manufacturers. This was achieved by training the model on cells from two manufacturers and testing it on cells from the third. This approach evaluates how well the model, trained on cells from specific manufacturers, can predict the performance of cells from a different manufacturer. Since manufacturing processes introduce variability in electrode design, assembly, and material quality, this validation is critical for ensuring the model’s robustness across different production sources.
@@
For completeness, a standard randomized 5-fold cross-validation was also performed, and the results are presented in the Appendix. However, the results presented in this manuscript focus on the validation strategies discussed above which are tailored to the specific challenges of generalization across operating conditions and manufacturers.
@@ 4.1 Cross voltage and temperature performance
To establish a performance baseline for the model across different operational conditions, an established and widely applied feature set is first examined: the original six-feature $\Delta$Q(V) set from (Severson et al., 2019). This feature set has been extensively used in prior work for cycle life prediction and serves as a benchmark for comparison with other feature sets. Fig. 3a shows the parity plot comparing predicted and actual EOL85 values using the leave one triplicate out cross-validation strategy, which was discussed in the previous section. In all regression tasks, error is evaluated using the mean MAE of all cross-validation (CV) validation set predictions.
Figure 3: Predicted vs. actual cycles to EOL for different feature sets from an elastic net regressor across varying conditions. Panels (a) and (e) depict predictions using the feature set from (Severson et al., 2019), while panels (b) and (f) show predictions using a downselected set of 20 features derived from novel DCIR-based $\Delta Q(V)$ data, comparing cycles 102 and 2. Panels (c) and (g) demonstrate results from step-wise backward sequential feature selection, reducing all features to 20 for validation across different testing conditions (top row) and to 10 features for validation across different manufacturers (bottom row), respectively. (d) $R^{2}$ scores for different feature sets, corresponding to the predicted number of cycles to reach EOL80, EOL85, and EOL90.
This baseline model is directly compared with an equivalent model trained on an augmented feature set, which includes univariate time-series features computed from DCIR pulses. The expanded feature set consists of 60 attributes, significantly increasing the dimensionality and predictive power of the model. The inclusion of augmented DCIR $\Delta$Q(V) features captures intricate degradation patterns that are not discernible using the standard feature set alone. As demonstrated in Fig.3b this approach results in more accurate predictions, with a notable improvement in MAE of 161 cycles. Further, using an amalgamation of the downselected features from the entire feature set, including resistance values from the 10 DCIR pulses, conventional $\Delta$Q(V) univariate features, and linear fit parameters ($p_{1},p_{2}$) further enhances the EOL prediction (Fig.3c). Predictably, models exhibit lower errors at higher SOH thresholds, reflecting the reduced complexity of predicting less degraded states, as presented in Table 3.
The errors and scores reported in Table 3 and Fig. 3a-d are based on the custom CV strategy outlined in the previous section. For reference, using a more conventional 5-fold CV strategy, as detailed in the SI, results in even better performance. However, given the uniqueness of the dataset, which includes cell triplicates, the custom CV strategy is argued to better showcase the model’s generalizability across testing conditions.
SOH (%) $\Delta Q(V)$ non-DCIR $\Delta Q_{\text{DCIR}}(V)$ All features
80 424.8 425.9 294.3 179.6
85 390.6 319.8 160.6 113.4
90 197.1 142.1 107.4 72.6
Table 3: MAE for different feature combinations for EOL defined at different SOH thresholds. Features based on DCIR measurements provide significant improvements in prediction accuracy compared to non-DCIR features.
4.2 Cross manufacturer performance
Traditional cycle life prediction using $\Delta Q(V)_{101-1}$ features, though simple and widely used, fail to generalize across manufacturers, as shown in Fig. 3e with an MAE of 416 cycles. These features capture broad degradation trends but overlook critical indicators like internal resistance growth, which results from processes such as SEI formation and active mass loss due to electrode cracking and structural transformations. Additionally, variations in electrode design and manufacturing processes further limit the predictive power of $\Delta Q(V)$ alone.
The incorporation of early-cycle DCIR-based features, as illustrated in Fig. 3f with an MAE 265 cycles, substantially improves the model’s generalization across manufacturers. DCIR tracks the evolution of internal resistance across various states of charge. This provides a more comprehensive assessment of the cell’s internal electrochemical processes, which enables the model to capture nuanced changes in degradation behavior.
Furthermore, the combination of both DCIR and traditional $\Delta Q(V)$ features, as shown in Fig. 3g with an MAE of 141 cycles, results in a marked improvement in prediction accuracy. By leveraging both the broad degradation trends captured by $\Delta Q(V)$ and the detailed internal resistance information provided by DCIR, the model can more accurately predict state-of-health across a wide range of cell designs and manufacturing processes. This hybrid feature set is particularly effective for bridging the gap between cells with different electrode materials, as evidenced by the model’s ability to predict the EOL of LG graphite-only cells based solely on training data from silicon-graphite cells produced by Samsung and BAK, as well as the accurate prediction of Samsung’s NCA-based cells based the NMC cells by LG and BAK.
@@ 5 Electrochemical Indicators and Their Role in Generalization
Variations in manufacturing practices, such as electrode composition, coating uniformity, and assembly precision, introduce significant complexity into cycle life prediction. These subtle differences manifest as changes in internal resistance, capacity fade, and overall cell degradation. Internal resistance, in particular, serves as a critical indicator of underlying degradation mechanisms, including solid electrolyte interphase (SEI) growth, particle fracture, and lithium plating. These degradation processes influence the electrochemical behavior of the cell, which in turn impacts its performance and lifetime.
@@
Figure 4: Feature-based drivers of model performance. SHAP values (in cycles) demonstrating the contribution of various features to the prediction of EOL85 across (a) operating conditions and (b) manufacturers. Positive SHAP values indicate a positive contribution to longer cycle life, while negative values imply a reduction in predicted cycle life. The color gradient corresponds to the magnitude of feature values, where darker colors indicate higher values. (c) Voltage difference plotted against $\Delta Q_{DCIR,102-2}^{(0)}$ (Ah) for cells from three different manufacturers (LG - red, Samsung - blue, BAK - green). Darker colors correspond to a greater number of cycles to EOL.
@@
As shown in Fig. 4, resistance-based features ($R_{102}$, $R_{2}$) and other impedance-related indicators play a critical role in predicting cycle life, contributing substantially to model performance under various operating conditions (Fig. 4a) and across different manufacturers (Fig. 4b). These features, which track resistance behavior at different states of charge, likely capture degradation mechanisms, such as solid electrolyte interphase (SEI) growth and lithium plating, that can manifest across different chemistries and manufacturing processes, though the specific behavior of these mechanisms may still be influenced by these factors. As illustrated in Fig. 4c, cells with greater number of cycles at EOL85, as indicated by the darker colors, exhibit smaller $\Delta Q_{DCIR,102-2}^{(0)}$. This suggests that impedance-based features can effectively differentiate degradation behaviors even when the underlying cell chemistries and manufacturing processes vary significantly.
@@
In contrast, capacity-based features primarily track shifts in capacity fade but may fail to capture early-stage degradation mechanisms such as internal resistance growth, which can occur well before significant capacity loss is observed. This limitation becomes particularly evident when attempting to generalize models across different manufacturers, as subtle variations in manufacturing processes (e.g., electrode coating uniformity or electrolyte formulation) can result in different onset points and rates of degradation. The ability of impedance-based features to generalize across these variations highlights their importance in achieving reliable lifetime predictions.
@@
Overall, the ability of DCIR-based features to reflect internal resistance growth, irrespective of specific manufacturing differences, contributes significantly to the model’s robustness in generalizing across a wide range of cell designs. These features effectively capture the microscopic and macroscopic degradation processes, ensuring a more reliable prediction of battery life despite variations in cell chemistry or manufacturing techniques.
@@ 6 Conclusion
This work demonstrates the efficacy of incorporating internal-impedance-based features from pulse measurements into data-driven models for the prediction of Li-ion battery state-of-health, addressing limitations associated with traditional capacity-based features. While these conventional features capture broad degradation trends, they fail to generalize effectively across different manufacturers due to variations in cell design, electrode composition, and manufacturing processes. The inclusion of DCIR data, particularly in early cycles, allows for the detection of internal resistance growth, which is closely tied to key degradation mechanisms. Results presented in this work show that this integrated approach significantly improves prediction accuracy and generalizes well across different manufacturers and chemistries. Moreover, this methodology has the potential to allow manufacturers to adapt their battery compositions and leverage pre-existing data for predictive modeling without the need to build entirely new datasets, facilitating rapid innovation and optimization of new cell designs. Incorporating DCIR-based features into predictive models could thus provide a more robust and efficient framework for EOL estimation, advancing battery management systems and their performance in diverse real-world applications.
7 Data Availability and Future Revisions
The data supporting the findings of this study will be made publicly available in the coming weeks. Additionally, minor edits and further content updates are expected as the manuscript undergoes refinement based on ongoing review and feedback.

### v2

@@ 3.2 Cross-validation strategies
First, since the dataset is composed of cells grouped into triplicates, where each triplicate is a set of three cells cycled under identical conditions (i.e., a specific combination of manufacturer, temperature, and voltage range), a leave one triplicate out (LOTO) cross-validation strategy was adopted to assess the model’s ability to generalize across unique samples. In this approach, the model was trained on all but one triplicate, and the remaining triplicates were used for testing. This process was repeated until all triplicates were used as test sets, thus evaluating the model across a diverse range of operating conditions and cell types. By excluding entire triplicates during each fold, this method tests the model’s ability to predict cycle life for a particular cell type and unseen operating condition. This strategy maximizes the utility of the dataset and provides an assessment of the model’s generalization capabilities across temperature, voltage, and manufacturer combinations.
Second, the LOTO strategy was extended to leave one operating condition out (LOOCO) in which folds are created by omitting from model training all cells tested under a particular combination of temperature and voltage range. This approach assesses how well the feature set generalizes to unseen operating conditions.
Third, leave one manufacturer out (LOMO) cross-validation was performed to assess the model’s ability to generalize across different cell types. This was achieved by training the model on cells from two manufacturers and testing it on cells from the third. It is important to emphasize that with this approach, the models were tested on unseen cell types across the various temperature and voltage range combinations. This approach evaluates the capacity of the feature set to generalize to new cell types. It is important to note that the cell types in this dataset have in common Ni-based cathodes and Gr-based anodes, thus the LOMO strategy does not validate across other chemistry families, such as LFP, for example. However, it is worth re-iterating that the three cell types considered here were each distinct in their exact cathode-anode compositions (e.g, NCA and NMC, or Gr-Si and Gr). The LOMO strategy is an important consideration for assessing model performance because manufacturing processes introduce variability in electrode design, assembly, and material quality, thus ensuring robustness across different production sources.
@@
For completeness, a standard randomized 5-fold cross-validation was also performed, and the results are presented in the Appendix. In general, predictions are better using 5-fold cross-validation because the models were trained on 1 or more from a triplicate set, while predicting on a cell from the same set. The results presented in this manuscript focus on the cross-validation strategies discussed above which are tailored to the specific challenges of generalization across operating conditions and manufacturers.
@@ 4 Results and Discussion
To establish a performance baseline for model performance across operating conditions and cell types, predictions are compared against features computed from the differences in voltage vs capacity profiles between cycles 101 and 1 ($\Delta Q_{101-1}(V)$, described in section 3.1) and using the leave one triplicate out cross-validation strategy (LOTO, described in section 3.2). These features have been well-studied in the literature Severson et al. (2019) thus provide a good reference point. In the following discussion, comparisons are made across feature sets and cross-validation strategies to assess model generalizabilty.
Fig. 3a-c shows parity plots comparing predicted and actual EOL85 with the LOTO cross-validation strategy. Fig. 3a,b,c respectively show predictions made with models trained on features computed from voltage profile differences ($\Delta Q_{101-1}(V)$), current pulse differences ($\Delta Q_{DCIR}(V)$), and a combination of all features that additionally consider direct-current internal resistance ($R_{DCIR}$) and linear fit parameters to the lifetime curves. In each case, the feature sets are down-selected to 20 features. The predictions systematically improve when current pulses are included in the feature selection. Fig. 4a shows the MAE averaged over all cross-validation folds for the LOTO cross-validation strategy and for three EOL thresholds (90%, 85%, and 80%). The green squares (EOL85) correspond to the average MAE for the parity plots of fig. 3a-c. The average MAE is reduced from more than 400 cycles to less than 200 cycles for the EOL80 predictions and from  400 cycles to nearly 100 cycles for the EOL85 predictions. The differences in model MAE for the various EOL conditions can be attributed to two factors: i) some cells did not reach the EOL80 or EOL85 threshold which means there were fewer cells in then training set, and ii) in earlier cycles the lifetime curves are highly linear, whereas in later life some of these curves exhibit non-linear behavior that makes predictions more challenging. Nevertheless, the inclusion of features from current pulses ($\Delta Q_{DCIR}(V)$ and $R_{DCIR}$) reduces the prediction error by a factor of 2 to 3.
Figure 3: Predicted vs. actual cycles to EOL for different feature sets from an elastic net regressor across varying conditions. Panels (a) (d) and (g) depict predictions using the feature set from Severson et al. (2019). Panels (b), (e) and (h) show predictions using a downselected set of features from DCIR current pulse differences at various states of charge between cycles 102 and 2. Panels (c), (f), and (i) show results for a down-selection of all features considered in this work. The top, middle, and bottom rows correspond to leave one triplicate out, leave one operating condition out, and leave one manufacturer out cross validation strategies, respectively.
Figure 4: Prediction Mean Absolute Error across feature sets, end-of-life thresholds, and cross-validation strategies. Mean Absolute Error (MAE) against feature set averaged over all cross-validation folds using (a) leave-one triplicate out (LOTO), (b) leave-one operating condition out (LOOCO), and (c) leave-one manufacturer out (LOMO) cross-validation strategies and for 90% (purple), 85% (green), and 80% (yellow) end-of-life thresholds.
@@ 4.1 Cross voltage and temperature performance
Fig. 3d-f depicts how well model predictions generalize to unseen temperatures and voltage ranges (LOOCO cross-validation described in section 3.2) across feature sets and using an EOL threshold of 85%. The same three sets of derived features as fig. 3a-c were used but down-selected to 12 features since each cross-validation fold contained fewer samples. Predictions improve significantly with the inclusion of features computed from current pulses; fig. 3e shows a noticeable tightening of the predicted vs actual EOL data to the ideal line (black-dotted) compared to fig. 3d. This particular comparison is interesting because in both scenarios univariate time-series features from differences in voltage vs capacity profiles are used, thus highlighting that current pulses spanning the available SOC range encode important degradation information in the electrochemical signature. To a greater degree, injecting phenomenological insight in the feature generation (fig. 3f) allows for simple feature construction from current pulses by computing an effective resistance ($R_{DCIR}$) for each pulse and improves the prediction error further. Section 5 elaborates on the importance of various features. Prediction errors are summarized in fig. 4b, showing the averaged MAE for the LOOCO strategy and the three EOL conditions. The green squares correspond to panels d-f in fig. 3. When attempting to predict EOL with an 85% capacity threshold (green squares), the standard $\Delta Q(V)$ features yield an average MAE of over 400 cycles, while the optimized feature set that leverages current pulses yields an MAE just over 100 cycles; nearly four-fold reduction in prediction error. This demonstrates that features based on current pulses capture voltage- and temperature-dependent degradation modes that are not discernible from single current voltage profiles.
4.2 Cross cell manufacturer performance
Fig. 3g,h,i demonstrate how features based on current pulses enable models to generalize well to new cell types by using the leave one manufacturer out (LOMO) cross-validation strategy presented in section 3.2. In each of the three panels, the same feature sets as fig. 3a,b,c were used but down-selected to 12 features because each cross-validation fold contained fewer samples compared to the LOTO strategy. Similarly to generalization to new operating conditions, the inclusion of univariate time-series features based on current pulses (fig. 3h) improves predictions drastically. The inclusion of features based on effective resistances ($R_{DCIR}$) further tightens the cycle life predictions to the actual values (fig. 3i). Fig. 4c summarizes predictions using the LOMO cross-validation strategy for the three EOL thresholds and three feature sets considered. The EOL85 data (green squares) correspond to the average MAE from fig. 3g,h,i. For cell cycle life degradation beyond 90%, the improvement in prediction error when current pulse-based features are included is accentuated; the averaged MAE for both EOL85 and EOL80 are reduced by more than half. Considering that these cycle life prediction errors correspond to cells types that were unseen in training (within each respective cross-validation fold), the inclusion of current pulses, and to a greater extent effective resistances, leads to models that encode information on degradation processes not unique to a single cell chemistry.
4.3 Model generalizability summary
The averaged MAE summarized in fig. 4 illustrates the importance of encoding the electrochemical information that spans the entire SOC range, with the inclusion of pulse voltage-capacity profiles, and further illustrates the importance of injecting phenomenology in feature generation, with the inclusion of effective resistances. Two points of interests are discussed with regards to the current work: I) the increase in feature set dimension with the inclusion of current pulses, and II) the prediction errors of the voltage-capacity profile features ($\Delta Q(V)$). I) after down-selection, the feature space increases dimension from 6 to 20 (LOTO), from 6 to 12 (LOOCO), and from 6 to 10 (LOMO), thus prediction errors are expected to decrease. However, this is counter-balanced by II) examining fig. 3a,d,g again, it is observed that $\Delta Q(V)$ features make predictions within a band of values in which short-lived cells are over-predicted and long-lived cells are under-predicted, not effectively capturing trends. It can thus be postulated that by including additional long-lived cells the MAE would get inflated. This is not a phenomena solely attributable to feature space dimension, rather it shows the importance of resolving features across the entire available SOC range.
Of particular importance is how close the the prediction errors are in fig. 4a,b,c for each EOL thresholds and all three cross-validation strategies when considering all four feature sets ($\Delta Q(V)$, $\Delta Q_{DCIR}(V)$, $R_{DCIR}$, $p_{0}$, $p_{1}$). It is recalled that since some cells degraded slowly, the number of samples available for training decreased as the EOL threshold decreased; fewer cells reached 80% SOH compared to 85%, and again compared to 90%. This could certainly contribute to the general increase in prediction error as the EOL threshold is reduced. Nevertheless, considering the varied degradation trends exhibited within this dataset (fig. 1), the accurate and consistent cycle life predictions made by models that leverage features from current pulses and phenomenological insight across each EOL condition suggests that different degradation modes are encoded by these features, allowing for accurate predictions even during rapid cell failure. For example, sudden cell failure is generally accompanied with a rapid increase in internal resistance, which makes evident why cycle life prediction models that use $\Delta Q(V)$ features fail to generalize across degradation modes: the detailed electrochemical signature is overlooked. Developing features from current pulses that span the available SOC range appears to capture detailed differences in degradation such as solid-electrolyte interphase (SEI) growth, active material loss due to particle cracking or structural transformations, and electrolyte salt and solvent depletion. For example, even within a single cell type, differences in cycling voltage can have a dramatic impact on degradation; a Si-containing anode (Sam50E, BakCG) can operate effectively as a graphite-only anode if the lower voltage limit is sufficiently high, mitigating Si-related degradation modes. Thus, varying operating conditions, cell chemistries, and manufacturing processes lead to differences in degradation modes, and the ability to predict cycle life across these factors provides evidence of the rich information encoded in features computed from current pulses that span the entire SOC range.
@@ 5 Electrochemical Indicators and Their Role in Generalization
Differences in battery cell manufacturing such as electrode composition, electrode slurry recipes, electrode coating thickness and uniformity, cell form factors, and assembly precision, introduce significant complexity for cycle life prediction. Subtle differences between cells can manifest during cycling and aging as changes in internal resistance, capacity fade, and overall cell degradation. Internal resistance, in particular, can serve as an indicator of underlying degradation mechanisms, including solid electrolyte interphase (SEI) growth, electrode particle fracture, and lithium plating. These degradation processes influence the electrochemical behavior of the cell, which in turn impacts its performance and lifetime. The following discussion discerns feature importance and offers connection to electrochemical phenomena.
@@
Figure 5: Feature-based drivers of model performance. SHAP values (in cycles) showing the contribution of various features to the prediction of EOL85 for (a) leave one operating condition out and (b) leave one manufacturer out cross-validation strategies. The features are ordered from top to bottom by relative importance for model prediction. Positive SHAP values indicate a positive contribution to longer predicted cycle life, while negative values imply a reduction in predicted cycle life. The color gradient corresponds to the magnitude of feature values, where darker colors indicate higher values. (c) Voltage difference plotted against $\Delta Q_{DCIR,102-2}^{(0)}$ for cells from three different manufacturers (LG - red, Samsung - blue, BAK - green). Darker colors correspond to a greater number of cycles to EOL.
@@
Fig. 5 shows that resistance-based features ($R^{(i)}_{102}$, $R^{(j)}_{2}$) play a critical role in predicting cycle life, contributing substantially to model performance under various operating conditions (Fig. 5a) and across different manufacturers (Fig. 5b). These features track changes in SOC-dependent cell resistance due to degradation mechanisms such as solid electrolyte interphase (SEI) growth, electrolyte oxidation, lithium plating, cathode particle surface reconstruction, and poor electrode connectivity that can manifest across different chemistries and manufacturing processes. It is important to realize that the individual SHAP values in Fig. 5a-b should not be directly extended to electrochemical analysis; models encode relationships between the 12 values to make predictions. Nevertheless, the relative importance of features does indicate that resistance-based features derived from current pulses provide rich information with generalizable predictive power. Fig. 5c illustrates that cells with a greater number of cycles at EOL85, indicated by the darker colors, exhibit smaller $\Delta Q_{DCIR,102-2}^{(0)}$. This suggests that features based on current pulses across the SOC range can effectively differentiate degradation behaviors even when the underlying cell chemistries and manufacturing processes vary significantly.
@@
In contrast, features based on differences in voltage-capacity profiles primarily track shifts in capacity fade but fail to capture early-stage degradation mechanisms that lead to internal resistance growth, which can occur well before significant capacity loss is observed. This limitation becomes particularly evident when attempting to generalize models across different manufacturers, as subtle variations in manufacturing processes (e.g., electrode coating uniformity or electrolyte formulation) can result in different onset points and rates of degradation. The ability of resistance-based features that span the available SOC range to generalize across varied conditions highlights their importance in achieving reliable lifetime predictions.
@@
Overall, the ability of DCIR features ($R^{(i)}_{DCIR}$) to reflect internal resistance growth, contributes significantly to the model’s robustness in generalizing across cell types and operating conditions. These features effectively capture early-onset degradation processes, ensuring more reliable battery lifetime predictions despite varied operating conditions, cell chemistries, and manufacturing processes.
@@ 6 Conclusion
This work demonstrates the efficacy of incorporating features from current pulse measurements into data-driven models for the prediction of Li-ion battery cycle life, addressing limitations associated with traditional voltage-capacity-based features. While these conventional features capture broad degradation trends, they fail to generalize effectively across different temperatures, voltage ranges, and cell manufacturers due to variations in cell design, electrode composition, and manufacturing processes. The inclusion of DCIR data, particularly in early cycles, allows for the detection of internal resistance growth, which is closely tied to key degradation mechanisms. Results presented in this work show that this integrated approach significantly improves prediction accuracy and generalizes well across different manufacturers and chemistries. Moreover, this methodology has the potential to allow manufacturers to adapt their battery compositions and leverage pre-existing data for predictive modeling without the need to build entirely new datasets, facilitating rapid innovation and optimization of new cell designs. Incorporating DCIR-based features into predictive models could thus provide a more robust and efficient framework for EOL estimation that can accelerate research and development and potentially provide early warnings for in-field applications.
7 Methods
7.1 Data generation
57 ready-for-use commercial Li-ions cells were obtained for this work. Included were: 21 Samsung INR21700-50E, 21 BAK N21700CG, and 15 LG INR18650-MH1 cells. Cell operating conditions included 3 voltage ranges (2.5 - 4.06 V, 2.5 - 4.2 V, and 3.3 - 4.2 V) and three temperatures (23, 40, and 60 °C). Tests at 60 °C were only performed within the 2.5 - 4.2 V range and the LG INR18650-MH1 cells were not tested within the 3.3 - 4.2 V range. Tests were carried out with triplicate cells (3 cells from the same manufacturer operated at the same conditions and with the same charge-discharge protocol). The experiment design thus consisted of 19 unique samples (combination of cell type, temperature, and voltage range). Five cells experienced early failure with internal current interruption and were omitted from the dataset and anlysis: 1 Samsung INR21700-50E (23 °C, 3.3 - 4.2 V), 1 BAK N21700CG (23 °C, 2.5 - 4.06 V), and 3 BAK N217000CG (60 °C, 2.5 - 4.2 V). Additionally, two LG INR18650-MH1 (25 °C, 2.5 - 4.06 V) cells never degraded to 90 % or lower and were ommitted from the dataset. Cell cycling was carried out on NEWARE BTS-4000 cyclers. For cells tested at 40 °C, the temperature was maintained using in-house customized industrial thermal chambers, with estimated maximum temperature fluctuations of 1 °C. For cells tested at 60 °C, the temperature was maintained with NOVONIX thermal chambers, with estimated maximum temperature fluctuations of less than 0.1 °C. Cells tested at 23 °C were in a temperature-controlled warehouse with estimated maximum temperature fluctuations of 1 °C. The charge-discharge sequence employed for all cells follows a 100 cycle pattern: C/10 charge, C/10 discharge, C/10 charge, C/10 discharge with 1C-equivalent pulses followed by a 15 minute rest (open-circuit) each 10 % SOC, 98 cycles at C/3 (constant-current charge with a constant-voltage hold at top-of-charge terminating when the current is less than C/20 rate equivalent, followed by constant current discharge). Sample protocol files have been included in the data repository listed in the Data Availability section.
7.2 Data processing
The files output by the cycler systems were processed and stored in the Apache Arrow Feather file format. Some quantities were computed during processing but the original data was preserved through the file store conversion. In the data repository quoted in the Data Availability section, only the Feather files are made available because they contain the original data in addition to all computed quantities. Aggregated quantities (per-cycle metrics) are stored in a separate Feather file. Data dictionaries are provided in the open-access data repository.
7.3 Feature generation
The present work focuses on systematically deriving features from the first 102 cycles, encompassing both full discharge curves and discrete voltage responses during DCIR current pulses. Following the methodology of  Severson et al. (2019), a structured interpolation scheme was employed for the capacity–voltage data to facilitate cycle-to-cycle comparisons and ultimately extract statistical descriptors that capture changes in the battery’s state of health.
Traditional Capacity-based features. For each cycle, the discharge capacity ($Q$) was measured as a function of voltage ($V$). To enable a consistent pointwise comparison between cycles, a voltage window was defined spanning the full discharge range for each cell (e.g., 2.5–4.2 V), over which linear interpolation of $Q(V)$ was performed. Each $Q(V)$ curve was sampled at a uniform voltage grid, ensuring that capacity data from different cycles could be directly subtracted at the same voltage points:
$Q_{n}(V)\;\longrightarrow\;\widetilde{Q}_{n}(v_{i}),$
where $\widetilde{Q}_{n}(v_{i})$ denotes the interpolated capacity at the $i$-th voltage point $v_{i}$.
The difference between two representative cycles $m$ and $n$ was then computed:
$\Delta Q_{m-n}(v_{i})\;=\;\widetilde{Q}_{m}(v_{i})\;-\;\widetilde{Q}_{n}(v_{i}).$
In this study, $\Delta Q_{101-1}(V)$ was selected to capture early-to-later cycle changes in capacity at the slower C/10 discharge rate. Following  Severson et al. (2019), a suite of univariate statistical descriptors was extracted from $\Delta Q(V)$, including the mean ($\mu$), minimum ($\min$), maximum ($\max$), variance ($\sigma^{2}$), skewness ($\gamma_{1}$), and kurtosis ($\gamma_{2}$). These descriptors summarize the shape and distributional changes in the discharge curves over time, reflecting diverse degradation mechanisms (e.g., loss of active material, shifts in electrode kinetics).
DCIR-based features. During the checkup cycles, DCIR measurements were performed by applying short current pulses at specific state-of-charge (SOC) levels and recording the resultant voltage drop. The effective internal resistance, $R_{\text{DCIR}}=\Delta V/\Delta I$, was recorded for each of the 10 SOC checkpoints in both an early cycle (cycle 2) and a later cycle (cycle 102). These measured resistances were used directly as features ($R_{\text{DCIR},k}$, where $k\in\{1,2,\ldots,10\}$).
In addition to the resistance measurements, the current pulses in each DCIR checkup define partial discharge segments, each providing a localized $Q(V)$ relationship in a narrower SOC range. By concatenating the voltage and capacity data within each of these partial segments and using the same interpolation approach described above, partial $Q(V)$ curves were formed for each pulse number $k$. The difference between the interpolated partial capacity curves in cycles 102 and 2, $\Delta Q_{\text{DCIR},102-2}^{(k)}(V)$, was then computed for each SOC checkpoint. The same univariate statistical descriptors (mean, minimum, maximum, variance, skewness, kurtosis) were extracted for each $\Delta Q_{\text{DCIR}}^{(k)}(V)$. This yielded 60 additional features (6 statistics $\times$ 10 pulses), collectively referred to as $\Delta Q_{\text{DCIR}}^{(k)}$ features. These metrics capture localized capacity evolution at various SOC levels, enhancing sensitivity to distinct degradation processes that may manifest in specific regions of the voltage curve.
Lifetime Curve Features. To capture broad trends in capacity fade, a linear fit of the discharge capacity vs. cycle number was applied from an early to an intermediate stage of cycling (e.g., cycles 30–99). Let $Q_{\mathrm{dch}}(c)$ denote the capacity at cycle $c$. The linear model was:
$Q_{\mathrm{dch}}(c)\;=\;p_{1}\,c+p_{2},$
yielding slope ($p_{1}$) and intercept ($p_{2}$) as features. The slope provides an estimate of the rate of capacity loss during the chosen interval, while the intercept reflects the baseline capacity at the beginning of the fit window. Although higher-order polynomial fits were considered, they did not yield significant performance improvements in subsequent predictive modeling.
This comprehensive feature set targets both global and localized aspects of cell aging, leveraging repeated high-resolution measurements of voltage, capacity, and resistance. By combining the established framework of Severson et al. Severson et al. (2019) with supplemental partial-discharge descriptors from DCIR checkups, richer insights into the diverse pathways of battery degradation are enabled.
7.4 Model selection and training
To simultaneously perform model fitting and feature selection, a regularization-based linear regression approach was adopted. A linear model of the form
$\hat{y}=\hat{w}X$
was employed, where $\hat{y}_{i}$ denotes the predicted cycle life for cell $i$, $X_{i}$ is the feature vector for cell $i$, and $\hat{w}$ is the coefficient vector. The elastic net regression was chosen due to its effectiveness in handling highly correlated features and relatively small sample sizes, which is characteristic of battery lifetime datasets. The elastic net formulation optimizes:
$\hat{w}=\text{argmin}_{w}\left(\|y-Xw\|_{2}^{2}+\lambda\left[\alpha\|w\|_{1}+(1-\alpha)\|w\|_{2}^{2}\right]\right)$
where $\lambda$ is a non-negative regularization strength scalar, and $\alpha$ balances between the L1 norm ($\|w\|_{1}$) promoting sparsity and the L2 norm ($\|w\|_{2}^{2}$) promoting stability and handling collinearity.
Prior to model fitting, Sequential Feature Selection (SFS) was applied to systematically reduce dimensionality from the full initial feature set, which includes resistance-based DCIR features, voltage-capacity derived $\Delta Q(V)$ statistics, and linear fit parameters. The optimal number of features was tuned separately for each cross-validation (CV) scheme described above (leave one triplicate out, LOTO; leave one operating condition out, LOOCO; and leave one manufacturer out, LOMO), with an imposed maximum of 20 features for the LOTO CV and 14 features for the LOOCO and LOMO schemes to mitigate potential overfitting. Optimum number of features was determined to be 20 for LOTO and 12 for LOOCO and 10 for LOMO.
Following feature selection, Principal Component Analysis (PCA) was performed on the reduced feature set to capture variance efficiently and further reduce multi-collinearity among selected features. The elastic net model was subsequently trained using the principal components extracted from PCA.
Hyperparameter optimization of the elastic net (regularization strength $\lambda$ and mixing parameter $\alpha$) was conducted via nested cross-validation to avoid data leakage, using only training folds within each CV scheme. Mean absolute errors (MAE) were computed across the held-out test folds for each CV scheme to evaluate predictive accuracy and generalizability. This approach balances model interpretability and robustness, making optimal use of the rich dataset provided by DCIR-based cycling data and ensuring a fair and rigorous assessment of the model’s ability to generalize across various manufacturers and cycling conditions.
To interpret feature importances in the original input space, we applied model-agnostic SHAP analysis to a custom-wrapped pipeline comprising sequential feature selection (SFS), principal component analysis (PCA), and ElasticNet regression. By exposing a .predict() interface on the full pipeline and supplying untransformed input data to a KernelExplainer, SHAP perturbations and attributions were computed with respect to the preprocessed (raw) feature set rather than post-PCA components.
8 Data Availability
The data generated and analyzed during this study, including raw cycling data, cell metadata, summary diagnostic metrics, and Neware cycling protocol files, are available at the Open Science Framework repository: [link redacted for journal review. Interested readers should contact authors for early access]